OpenStudy (anonymous):

Find four consecutive integers whose sum is 114.

OpenStudy (anonymous):

Let the four consecutive integers be x,x+1,x+2,x+3 given : x+(x+1)+(x+2)+(x+3)=114 4x=114-6 4x=108 x=108/4 = 27 so, the consecutive integers- x= 27 x+1 = 28 x+2= 29 x+3 = 30